A particle moves along the circular path of radius 15 cm with a constant angular acceleration of 4 rad/s 2 . If the initial angular speed of the particle is 5 rad/s, find angular displacement of the particle in 5 second.

  1. A particle moves along the circular path of radius 15 cm with a constant angular acceleration of 4 rad/s2. If the initial angular speed of the particle is 5 rad/s, find angular displacement of the particle in 5 second. 

Grade:

1 Answers

Rajdeep
5 years ago
HELLO THERE!
 
This question can be solved by using the second Equation of Motion:
 
Let’s analyze the given data:
Initial angular velocity (wo= 5 rad/s
Angular acceleration (\alpha= 4 rad/s2
We need to find the angular displacement (\theta)...
 
From second equation of motion, we have:
 
\theta = \omega_{o}t +\frac{1}{2}\alpha t^{2}
Put the value of wo\alpha and t (5 sec),
we get 
 
\theta = 75^{o}
 
THANKS!